OpenStudy (anonymous):

m is the midpoint of AB . find the coordinates of the third point when A(2,7) AND M(1,6) whats b?

OpenStudy (anonymous):

i know the answer is (0,5) but how do i do the work?

OpenStudy (anonymous):

Mid point of two pts A(x1,y1) and B(x2,y2)=(x1+x2)/2and (y1+y2)/2

OpenStudy (anonymous):

huh confuzed...

OpenStudy (anonymous):

the coordinates of the mid point is the average of the coordinates of the points

OpenStudy (mertsj):

\[\frac{x _{1}+x _{2}}{2}=1\] \[\frac{2+x _{2}}{2}=1\] \[2+x _{2}=2\] \[x _{2}=0\]

OpenStudy (mertsj):

Repeat the process for the y coordinate of the midpoint.

OpenStudy (mysesshou):

sorry, got ahead of myself. A(2,7) and M(1,6) 2-1 = 1, so that means it is one unit over. A is at 2, M is at 1, so B is at 0 7-6=1, so this one is also one unit over, A is at 7, M is at 6, B is at 5. B(0,5)

OpenStudy (mertsj):

\[\frac{y _{1}+y _{2}}{2}=6\] \[\frac{7+y _{2}}{2}=6\] \[7+y _{2}=12\] \[y _{2}=5\]

OpenStudy (mysesshou):

Mertsj's method is more proper. Especially for a more difficult problem when the answer is not given.

OpenStudy (mysesshou):

Mertsj's method is recommended for exams and HW.

OpenStudy (anonymous):

thanks alott

OpenStudy (mertsj):

yw